# Totally geodesic submanifolds of the homogeneous nearly Kähler 6-manifolds and their $\mathsf{G}_{2}$-cones

### v2

@@ 2.2.3. The almost product $\mathsf{S}^{3}\times\mathsf{S}^{3}$
This space is obtained via the Ledger-Obata construction from the group $\mathsf{SU}(2)$ (see for example [41]). We consider the product $\mathsf{G}=\mathsf{SU}(2)^{3}$ and the subgroup $\mathsf{K}=\Delta\mathsf{SU}(2)$ obtained by embedding $\mathsf{SU}(2)$ diagonally in $\mathsf{G}$. Then $\mathsf{G}$ acts on $M=\mathsf{S}^{3}\times\mathsf{S}^{3}=\mathsf{SU}(2)\times\mathsf{SU}(2)$ via the equation $(g,h,k)\cdot(x,y)=(gxk^{-1},hyk^{-1})$, and the isotropy subgroup at $o=(I,I)$ is $\mathsf{K}$, so we obtain that $M=\mathsf{G}/\mathsf{K}$. The Killing form of $\mathfrak{su}(2)$ is $\mathcal{B}(X,Y)=4\operatorname{tr}(XY)$, and the direct sum $\mathcal{B}\oplus\mathcal{B}\oplus\mathcal{B}$ is precisely the Killing form of $\mathfrak{g}$ whose inverse yields the standard homogeneous metric on $M$. Similarly to the previous cases, we will consider the renormalized metric given by $\langle(X_{1},X_{2},X_{3}),(Y_{1},Y_{2},Y_{3})\rangle=-\operatorname{tr}(X_{1}Y_{1})-\operatorname{tr}(X_{2}Y_{2})-\operatorname{tr}(X_{3}Y_{3})$. We denote by $\mathfrak{p}$ the orthogonal complement of $\mathfrak{k}=\Delta\mathfrak{su}(2)$ in $\mathfrak{g}$ and we consider the order three automorphism $\Theta\colon\mathsf{G}\to\mathsf{G}$ defined by $\Theta(g,h,k)=(h,k,g)$. The nearly Kähler complex structure $J$ is given as $J=\frac{2}{\sqrt{3}}\Theta_{*}+\frac{1}{\sqrt{3}}\operatorname{Id}_{\mathfrak{p}}$. Furthermore, if $L_{g}\colon\mathsf{S}^{3}\to\mathsf{S}^{3}$ denotes left multiplication by $g\in\mathsf{S}^{3}$, the almost product structure of $\mathsf{S}^{3}\times\mathsf{S}^{3}$ is the $\mathsf{G}$-invariant tensor $P$ of type $(1,1)$ defined by
@@
Consider the following basis of $\mathfrak{su}(2)$ in terms of elementary matrices:
@@
Consider the chain of inclusions $\Delta\mathsf{SU}(2)\subseteq\Delta_{1,3}\mathsf{SU}(2)\times\mathsf{SU}(2)_{2}\subseteq\mathsf{SU}(2)\times\mathsf{SU}(2)\times\mathsf{SU}(2)$. This gives rise to the homogeneous fibration $\mathsf{S}^{3}\hookrightarrow\mathsf{S}^{3}\times\mathsf{S}^{3}\to\mathsf{S}^{3}$ which is merely the projection on the first factor. Once again the fibers are totally geodesic. The vertical and horizontal subspaces at $o$ are given by
@@
The isometry group of $\mathsf{S}^{3}\times\mathsf{S}^{3}$ is $I(\mathsf{S}^{3}\times\mathsf{S}^{3})=(\mathsf{SU}(2)^{3}/\Delta\mathbb{Z}_{2})\rtimes\mathfrak{S}_{3}$ where $\mathfrak{S}_{3}$ denotes the symmetric group on three elements acting in the natural manner on $\mathsf{S}^{3}\times\mathsf{S}^{3}$, see [58, Lemma 3.3] for a proof.
@@ 3. Totally geodesic immersions in real analytic Riemannian manifolds
To conduct the global study of totally geodesic submanifolds, it is natural to assume that the ambient space is real analytic, since homogeneous spaces are examples of real analytic manifolds [38, Proposition 4.2]. As we will see, it is also necessary to allow self-interesctions in our submanifolds, leading to the study of totally geodesic immersions in analytic Riemannian manifolds.
@@
Let $M$ be a Riemannian manifold and $f\colon\Sigma\to M$ an isometric immersion of a connected Riemannian manifold $\Sigma$ to $M$. We say that $f$ is a totally geodesic immersion (and $\Sigma$ is a totally geodesic submanifold of $M$) if for every geodesic $\gamma$ of $\Sigma$ the composition $f\circ\gamma$ is also a geodesic of $M$. This condition is equivalent to the vanishing of the second fundamental form $\mathbb{II}$ of $\Sigma$. We refer the reader to [28, Appendix A] and [6, Section 10.3] for more details on totally geodesic immersions. A similar treatment can be found in [57].
@@
When studying general totally geodesic immersions, certain redundancies arise that need to be avoided. The first one is due to reparametrizations of the domain, which is handled by introducing the notion of equivalence. We say that two totally geodesic immersions $f_{i}\colon\Sigma_{i}\to M$ are equivalent if there exists an isometry $h\colon\Sigma_{1}\to\Sigma_{2}$ such that $f_{1}=f_{2}\circ h$. The second redundancy occurs when considering surjective local isometries. For instance, a totally geodesic embedding of $\mathbb{R}\mathsf{P}^{2}$ in $\mathbb{R}\mathsf{P}^{3}$ yields the same information as its composition $\mathsf{S}^{2}\to\mathbb{R}\mathsf{P}^{3}$ with the universal cover $\mathsf{S}^{2}\to\mathbb{R}\mathsf{P}^{2}$, though the latter is not an embedding. We can deal with this issue by working exclusively with compatible immersions. Let $f\colon\Sigma^{k}\to M$ be totally geodesic, and consider the Grassmann bundle $\mathsf{G}_{k}(TM)$ of $k$-planes in $TM$. Then $f$ induces a smooth map $\widetilde{f}\colon\Sigma\to\mathsf{G}_{k}(TM)$ by letting $\widetilde{f}(p)=(f(p),f_{*p}(T_{p}\Sigma))$ (for ease of notation, we will simply write $\widetilde{f}(p)=f_{*p}(T_{p}\Sigma)$). The map $f$ is said to be compatible if $\widetilde{f}$ is injective (this is the case, for example, if $f$ is injective). In other words, we allow $\Sigma$ to have self-intersections in $M$, but at each point of self-intersection the corresponding tangent spaces must be different. It turns out that any totally geodesic immersion can be factored through a compatible one [6, p. 272], so imposing this condition does not cause us to lose information. We note that in general, given compatible totally geodesic immersions $f_{1}\colon\Sigma_{1}\to\Sigma_{2}$ and $f_{2}\colon\Sigma_{2}\to M$, the composition $f_{2}\circ f_{1}$ may not be compatible. A counterexample will be provided in Remark 5.4.
@@
One sees that a compatible immersion $f\colon\Sigma\to M$ is determined up to equivalence by the image $\widetilde{f}(\Sigma)\subseteq\mathsf{G}_{k}(TM)$. Given two compatible totally geodesic immersions $f_{i}\colon\Sigma_{i}\to M$, we say that $f_{2}\colon\Sigma_{2}\to M$ extends $f_{1}\colon\Sigma_{1}\to M$ if there exists an injective local isometry $j\colon\Sigma_{1}\to\Sigma_{2}$ satisfying $f_{1}=f_{2}\circ j$, or equivalently, if $\widetilde{f}_{1}(\Sigma_{1})\subseteq\widetilde{f}_{2}(\Sigma_{2})$. This defines a partial ordering in the set of equivalence classes of compatible totally geodesic immersions. By [6, Proposition 10.3.2], every compatible totally geodesic immersion can be extended to an inextendable one (that is, an immersion which is maximal with respect to the partial ordering), and this extension is unique up to equivalence (we say in this case that the submanifold is inextendable [^1] Note that Heintze, Liu and Olmos [28] refer to these submanifolds as maximal totally geodesic submanifolds, whereas we prefer to reserve this term for another concept.). We thus aim to classify inextendable totally geodesic submanifolds up to congruence.
@@
Let $M$ be a real analytic Riemannian manifold, $p\in M$ and $V\subseteq T_{p}M$ a vector subspace. There exists a totally geodesic submanifold $\Sigma$ of $M$ whose tangent space at $p$ is $V$ if and only if $(\nabla^{k}R)_{p}$ preserves $V$ for all $k\geq 0$.
@@
For the sake of convenience, we will refer to the totally geodesic submanifold $f\colon\Sigma\to M$ as $\Sigma$ during the rest of this paper unless we need to specify the immersion. In most cases, the totally geodesic immersions that will appear are inclusions of embedded submanifolds $\Sigma\subseteq M$, so in this context there is no ambiguity in omitting the immersion.
@@
If $V\subseteq T_{p}M$ is a totally geodesic subspace and $X\in T_{p}M$ is any vector, it follows from Theorem 3.1 that $V$ is an invariant subspace for $R_{X}$ and $C_{X}$. In particular, if $T_{p}M=(T_{p}M)_{\lambda_{1}}\oplus\dots\oplus(T_{p}M)_{\lambda_{r}}$ is the eigenspace decomposition of $T_{p}M$ with respect to $R_{X}$, we have $V=(V\cap(T_{p}M)_{\lambda_{1}})\oplus\dots\oplus(V\cap(T_{p}M)_{\lambda_{r}})$. The same argument can be applied for $C_{X}$.
@@
In this paper we focus on real analytic manifolds which are not necessarily complete (see Section 6). Because the exponential map of a real analytic Riemannian manifold is also real analytic, one sees that for any compatible totally geodesic immersion $f\colon\Sigma\to M$, the domain $\Sigma$ is also a real analytic manifold and $f$ is a real analytic map. We will provide a criterion that allows us to determine whether a totally geodesic submanifold $\Sigma$ of $M$ is inextendable in terms of the maximal geodesics of $\Sigma$ and $M$. The key observation that makes use of real analyticity is the following: if $p\in M$ and $V\subseteq T_{p}M$ is a totally geodesic subspace, then for every geodesic $\gamma\colon I\subseteq\mathbb{R}\to M$ satisfying $\gamma(0)=p$ and $\gamma^{\prime}(0)\in V$, the parallel translates $\mathcal{P}_{0,t}^{\gamma}V$ are also totally geodesic subspaces of $T_{\gamma(t)}M$ for all $t\in I$ (see for example the proof in [57, Corollary 3.7]).
@@
Firstly, suppose that (i) holds. If (ii) is not satisfied, we can find a geodesic $\gamma\colon[0,1)\to\Sigma$ with $\gamma(0)=x$ that is not extendable to the right while the composition $f\circ\gamma$ admits an extension $\sigma\colon[0,1]\to M$. We consider $q=\sigma(1)$. The subspace $W=\mathcal{P}_{0,1}^{\sigma}V\subseteq T_{q}M$, where $V=\widetilde{f}(x)$ is therefore totally geodesic.
We show that $W\notin\widetilde{f}(\Sigma)$. Indeed, suppose that for a certain $y\in\Sigma$ we have $W=\widetilde{f}(y)$. Since $-\sigma^{\prime}(1)=-\mathcal{P}_{0,1}^{\sigma}\sigma^{\prime}(0)\in W=f_{*y}(T_{y}\Sigma)$, there exists a geodesic $\alpha\colon[0,\delta)\to\Sigma$ satisfying $\alpha(0)=y$ and $\alpha^{\prime}(0)=-(f_{*y})^{-1}\sigma^{\prime}(1)$. Suppose without loss of generality that $\delta<1$. The composition $f\circ\alpha$ satisfies $f(\alpha(0))=q$ and $(f\circ\alpha)^{\prime}(0)=-\sigma^{\prime}(1)$, so $f(\alpha(t))=\sigma(1-t)$ for $t\in[0,\delta)$. Furthermore, we have
@@
so using that $\widetilde{f}$ is injective we see that $\alpha(t)=\gamma(1-t)$ for all $t\in[0,\delta)$. Because $\alpha$ is continuous at 00, we see that the limit $\lim_{t\to 1^{-}}\gamma(t)$ exists and coincides with $y$, but this contradicts the fact that $\gamma$ is not extendable to the right. We deduce that $W$ is not in the image of $\widetilde{f}$.
@@
Thus, as a consequence of the Hopf-Rinow theorem and Proposition 3.4, one obtains the following corollary which generalizes a result by Hermann [29] to the case of non-complete ambient manifolds.

### v3

@@ 2.2.3. The almost product $\mathsf{S}^{3}\times\mathsf{S}^{3}$
This space is obtained via the Ledger–Obata construction from the group $\mathsf{SU}(2)$ (see for example [40]). We consider the product $\mathsf{G}=\mathsf{SU}(2)^{3}$ and the subgroup $\mathsf{K}=\Delta\mathsf{SU}(2)$ obtained by embedding $\mathsf{SU}(2)$ diagonally in $\mathsf{G}$. Then $\mathsf{G}$ acts on $M=\mathsf{S}^{3}\times\mathsf{S}^{3}=\mathsf{SU}(2)\times\mathsf{SU}(2)$ via the equation $(g,h,k)\cdot(x,y)=(gxk^{-1},hyk^{-1})$, and the isotropy subgroup at $o=(I,I)$ is $\mathsf{K}$, so we obtain that $M=\mathsf{G}/\mathsf{K}$. The Killing form of $\mathfrak{su}(2)$ is $\mathcal{B}(X,Y)=4\operatorname{tr}(XY)$, and the direct sum $\mathcal{B}\oplus\mathcal{B}\oplus\mathcal{B}$ is precisely the Killing form of $\mathfrak{g}$ whose opposite yields the standard homogeneous metric on $M$. Similarly to the previous cases, we consider the renormalized metric given by $\langle(X_{1},X_{2},X_{3}),(Y_{1},Y_{2},Y_{3})\rangle=-\operatorname{tr}(X_{1}Y_{1})-\operatorname{tr}(X_{2}Y_{2})-\operatorname{tr}(X_{3}Y_{3})$. We denote by $\mathfrak{p}$ the orthogonal complement of $\mathfrak{k}=\Delta\mathfrak{su}(2)$ in $\mathfrak{g}$ and we consider the order three automorphism $\Theta\colon\mathsf{G}\to\mathsf{G}$ defined by $\Theta(g,h,k)=(h,k,g)$. The nearly Kähler complex structure $J$ is given as $J=\frac{2}{\sqrt{3}}\Theta_{*}+\frac{1}{\sqrt{3}}\operatorname{Id}_{\mathfrak{p}}$. Furthermore, if $L_{g}\colon\mathsf{S}^{3}\to\mathsf{S}^{3}$ denotes left multiplication by $g\in\mathsf{S}^{3}$, the almost product structure of $\mathsf{S}^{3}\times\mathsf{S}^{3}$ is the $\mathsf{G}$-invariant tensor $P$ of type $(1,1)$ defined by
@@
We choose the following basis of $\mathfrak{su}(2)$ in terms of elementary matrices:
@@
Consider the chain of inclusions $\Delta\mathsf{SU}(2)\subseteq\Delta_{1,3}\mathsf{SU}(2)\times\mathsf{SU}(2)_{2}\subseteq\mathsf{SU}(2)\times\mathsf{SU}(2)\times\mathsf{SU}(2)$. This gives rise to the homogeneous fibration $\mathsf{S}^{3}\hookrightarrow\mathsf{S}^{3}\times\mathsf{S}^{3}\to\mathsf{S}^{3}$ which is merely the projection onto the first factor. Once again the fibers are totally geodesic. The vertical and horizontal subspaces at $o$ are given by
@@
The isometry group of $\mathsf{S}^{3}\times\mathsf{S}^{3}$ is $I(\mathsf{S}^{3}\times\mathsf{S}^{3})=(\mathsf{SU}(2)^{3}/\Delta\mathbb{Z}_{2})\rtimes\mathfrak{S}_{3}$ where $\mathfrak{S}_{3}$ denotes the symmetric group on three elements acting in the natural manner on $\mathsf{S}^{3}\times\mathsf{S}^{3}$, see [57, Lemma 3.3] for a proof.
@@ 3. Totally geodesic immersions in real analytic Riemannian manifolds
To conduct the global study of totally geodesic submanifolds, it is natural to assume that the ambient space is real analytic, since homogeneous spaces are examples of real analytic manifolds [37, Proposition 4.2]. As we will see, it is also necessary to allow self-intersections in our submanifolds, leading to the study of totally geodesic immersions in analytic Riemannian manifolds.
@@
Let $M$ be a Riemannian manifold and $f\colon\Sigma\to M$ an isometric immersion of a connected Riemannian manifold $\Sigma$ to $M$. We say that $f$ is a totally geodesic immersion (and $\Sigma$ is a totally geodesic submanifold of $M$) if for every geodesic $\gamma$ of $\Sigma$ the composition $f\circ\gamma$ is also a geodesic of $M$. This condition is equivalent to the vanishing of the second fundamental form $\mathbb{II}$ of $\Sigma$. We refer the reader to [27, Appendix A] and [6, Section 10.3] for more details on totally geodesic immersions. A similar treatment can be found in [56].
@@
When studying general totally geodesic immersions, certain redundancies arise that need to be avoided. The first one is due to reparametrizations of the domain, which is handled by introducing the notion of equivalence. We say that two totally geodesic immersions $f_{i}\colon\Sigma_{i}\to M$ are equivalent if there exists an isometry $h\colon\Sigma_{1}\to\Sigma_{2}$ such that $f_{1}=f_{2}\circ h$. The second redundancy occurs when considering surjective local isometries. For instance, a totally geodesic embedding of $\mathbb{R}\mathsf{P}^{2}$ in $\mathbb{R}\mathsf{P}^{3}$ yields the same information as its composition $\mathsf{S}^{2}\to\mathbb{R}\mathsf{P}^{3}$ with the universal cover $\mathsf{S}^{2}\to\mathbb{R}\mathsf{P}^{2}$, though the latter is not an embedding. We can deal with this issue by working exclusively with compatible immersions. Let $f\colon\Sigma^{k}\to M$ be totally geodesic, and consider the Grassmann bundle $\mathsf{G}_{k}(TM)$ of $k$-planes in $TM$. Then $f$ induces a smooth map $\widetilde{f}\colon\Sigma\to\mathsf{G}_{k}(TM)$ by letting $\widetilde{f}(p)=(f(p),f_{*p}(T_{p}\Sigma))$ (for ease of notation, we simply write $\widetilde{f}(p)=f_{*p}(T_{p}\Sigma)$). The map $f$ is said to be compatible if $\widetilde{f}$ is injective (this is the case, for example, if $f$ is injective). In other words, we allow $\Sigma$ to have self-intersections in $M$, but at each point of self-intersection the corresponding tangent spaces must be different. It turns out that any totally geodesic immersion can be factored through a compatible one [6, p. 272], so imposing this condition does not cause us to lose information. We note that in general, given compatible totally geodesic immersions $f_{1}\colon\Sigma_{1}\to\Sigma_{2}$ and $f_{2}\colon\Sigma_{2}\to M$, the composition $f_{2}\circ f_{1}$ may not be compatible. A counterexample will be provided in Remark 5.4.
@@
One sees that a compatible immersion $f\colon\Sigma\to M$ is determined up to equivalence by the image $\widetilde{f}(\Sigma)\subseteq\mathsf{G}_{k}(TM)$. Given two compatible totally geodesic immersions $f_{i}\colon\Sigma_{i}\to M$, we say that $f_{2}\colon\Sigma_{2}\to M$ extends $f_{1}\colon\Sigma_{1}\to M$ if there exists an injective local isometry $j\colon\Sigma_{1}\to\Sigma_{2}$ satisfying $f_{1}=f_{2}\circ j$, or equivalently, if $\widetilde{f}_{1}(\Sigma_{1})\subseteq\widetilde{f}_{2}(\Sigma_{2})$. This defines a partial ordering on any set of equivalence classes of compatible totally geodesic immersions. By [6, Proposition 10.3.2], every compatible totally geodesic immersion can be extended to an inextendable one (that is, an immersion which is maximal with respect to this partial ordering), and this extension is unique up to equivalence (we say in this case that the submanifold is inextendable). Note that Heintze, Liu and Olmos [27] refer to these submanifolds as maximal totally geodesic submanifolds, whereas we prefer to reserve this term for another concept. We thus aim to classify inextendable totally geodesic submanifolds up to congruence.
@@
Let $M$ be a real analytic Riemannian manifold, $p\in M$ and $V\subseteq T_{p}M$ a vector subspace. Then there exists a totally geodesic submanifold $\Sigma$ of $M$ whose tangent space at $p$ is $V$ if and only if $(\nabla^{k}R)_{p}$ preserves $V$ for all $k\geq 0$.
@@
For the sake of convenience, we refer to the totally geodesic submanifold $f\colon\Sigma\to M$ as $\Sigma$ during the rest of this paper unless we need to specify the immersion. In most cases, the totally geodesic immersions that will appear are inclusions of embedded submanifolds $\Sigma\subseteq M$, so in this context there is no ambiguity in omitting the immersion.
@@
If $V\subseteq T_{p}M$ is a totally geodesic subspace and $X\in T_{p}M$ is any vector, it follows from Theorem 3.1 that $V$ is an invariant subspace for $R_{X}$ and $C_{X}$. In particular, if $T_{p}M=(T_{p}M)_{\lambda_{1}}\oplus\dots\oplus(T_{p}M)_{\lambda_{r}}$ is the eigenspace decomposition of $T_{p}M$ with respect to $R_{X}$, we have $V=(V\cap(T_{p}M)_{\lambda_{1}})\oplus\dots\oplus(V\cap(T_{p}M)_{\lambda_{r}})$. The same argument can be applied for the operator $C_{X}$.
@@
In this paper we focus on real analytic manifolds which are not necessarily complete (see Section 6). Because the exponential map of a real analytic Riemannian manifold is also real analytic, one sees that for any compatible totally geodesic immersion $f\colon\Sigma\to M$, the domain $\Sigma$ is also a real analytic manifold and $f$ is a real analytic map.
We now provide a criterion that allows us to determine whether a totally geodesic submanifold $\Sigma$ of $M$ is inextendable in terms of the maximal geodesics of $\Sigma$ and $M$. The key observation that makes use of real analyticity is the following: if $p\in M$ and $V\subseteq T_{p}M$ is a totally geodesic subspace, then for every geodesic $\gamma\colon I\subseteq\mathbb{R}\to M$ satisfying $\gamma(0)=p$ and $\gamma^{\prime}(0)\in V$, the parallel translates $\mathcal{P}_{0,t}^{\gamma}V$ are also totally geodesic subspaces of $T_{\gamma(t)}M$ for all $t\in I$ (see for example the proof in [56, Corollary 3.7]).
@@
Firstly, suppose that (i) holds. If (ii) is not satisfied, then there exists a geodesic $\gamma\colon[0,1)\to\Sigma$ with $\gamma(0)=x$ that is not extendable to the right, while the composition $f\circ\gamma$ admits an extension $\sigma\colon[0,1]\to M$. We consider $q=\sigma(1)$. The subspace $W=\mathcal{P}_{0,1}^{\sigma}V\subseteq T_{q}M$, where $V=\widetilde{f}(x)$, is therefore totally geodesic.
We show that $W\notin\widetilde{f}(\Sigma)$. Indeed, suppose that for a certain $y\in\Sigma$ we have $W=\widetilde{f}(y)$. Since $-\sigma^{\prime}(1)=-\mathcal{P}_{0,1}^{\sigma}\sigma^{\prime}(0)\in W=f_{*y}(T_{y}\Sigma)$, there exists a geodesic $\alpha\colon[0,\delta)\to\Sigma$ satisfying $\alpha(0)=y$ and $\alpha^{\prime}(0)=-(f_{*y})^{-1}\sigma^{\prime}(1)$. Suppose without loss of generality that $\delta<1$. The composition $f\circ\alpha$ satisfies $f(\alpha(0))=q$ and $(f\circ\alpha)^{\prime}(0)=-\sigma^{\prime}(1)$, so $f(\alpha(t))=\sigma(1-t)$ for all $t\in[0,\delta)$. Furthermore, we have
@@
so using that $\widetilde{f}$ is injective we see that $\alpha(t)=\gamma(1-t)$ for all $t\in[0,\delta)$. Because $\alpha$ is continuous at 00, we obtain that the limit $\lim_{t\to 1^{-}}\gamma(t)$ exists and coincides with $y$, but this contradicts the fact that $\gamma$ is not extendable to the right. We deduce that $W$ is not in the image of $\widetilde{f}$.
@@
Thus, as a consequence of the Hopf–Rinow theorem and Proposition 3.4, one obtains the following corollary which generalizes a result by Hermann [28] to the case of non-complete ambient manifolds.